The measures of the angles of a triangle are in the extended ratio 7:9:10. What are the measures of the angles of the triangle. Can someone help me with this??
Answer by richwmiller(17219)

angles of a triangle add up to 180
a=7x,
b=9x,
c=10x,
7x+9x+10x=180
26x=180
13x=90
x=90/13
7*90/13= angle a
9*90/13= angle b
10*90/13= angle c
a = 630/13, b = 810/13, c = 900/13, x = 90/13
a=48.462, b=62.308, c=69.231, x=6.9231